OpenStudy (anonymous):

Please help me. Write the slope-intercept form of the line that has a slope of 2 and intersects the line, 2x - 3y = 6 at x = 3

OpenStudy (anonymous):

2 is the slope and -6 is the intercept

OpenStudy (anonymous):

(y-0) / (x-3) = 2

OpenStudy (anonymous):

y = 2*x - 6

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

is that all I need to put in or i need to do more steps?

OpenStudy (anonymous):

put it all in

OpenStudy (anonymous):

I did is that all I need to put in ?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

but put

OpenStudy (anonymous):

(y-0)/(x-3)=2 first

OpenStudy (anonymous):

ok i did